[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Tue Sep 27, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Wed Sep 28 02:26:03 UTC 2016


HSU Shift Log for Tue Sep 27 2016 16:17:08 GMT-0400 (EDT)

Shift: H

Forecaster: D Brown

Technical Issues Summary:

ATCF issues: None

N-AWIPS/AWIPS2 issues: Rebooted N-AWIPS machine around noontime. Machines  
came back up okay, and rebooting issue appears to be resolved.

Other Techincal issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20160928/96898815/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list